OpenStudy (anonymous):

Pls help Find an equation of the normal line to the graph of xe^xy = 7 at the point where x = 7 :) thank you :)

OpenStudy (sirm3d):

is the equation like this? \[\Large xe^{(xy)}=7\]

OpenStudy (amistre64):

should we assume the y is a function of x and do an implicit derivative?

OpenStudy (amistre64):

once you know the tangent line, you can always flip it around to a normal

OpenStudy (anonymous):

yes @sirm3d

OpenStudy (anonymous):

implicit derivative please @amistre64

OpenStudy (amistre64):

i take it you might be confused about how to operate implicit stuff then?

OpenStudy (anonymous):

yes :(

OpenStudy (amistre64):

hmm, do you know how to work the chain rule, the product rule, and the e^x rule ?

OpenStudy (anonymous):

yup

OpenStudy (amistre64):

then thats all there is to this problem, is just the name of the variables that is messing with your mind :)

OpenStudy (amistre64):

x e^(xy) = 7; lets clean it up by using a and b ab = 7; this is just a product rule ab'+a'b = 0 ; since a=x in this case, and b=the e parts .... xb'+x' e^(xy) = 0 ; and x' = 1 in this case xb'+e^(xy) = 0 all thats left is to determine the b'

OpenStudy (amistre64):

lets define b as: e^(pq) just to make it general we know e^f(x) = f'(x) e^f(x) right?

OpenStudy (amistre64):

the book might have it as: e^u derives to u' e^u

OpenStudy (anonymous):

so it will wnd up as y=1n1/7 ?

OpenStudy (amistre64):

since b=e^(pq); then b' is (pq)' e^(pq) (pq)' is just the product rule again: pq' + p'q, maybe, i havent seen the future yet :)

OpenStudy (anonymous):

i followed what you did.. and i think i got that :) hehe

OpenStudy (anonymous):

how do i get the equation of a normal line ?

OpenStudy (amistre64):

good :) x e^(xy) x (xy'+y) e^(xy) + e^(xy) = 0; and x=7 7 (7y'+y) e^(7y) + e^(7y) = 0 e^(7y) (49y'+7y +1) = 0 is there a value for y that you might have missed? or maybe we solve that in the original. the key is getting to y' and defining the tangent line at the given point

OpenStudy (amistre64):

once we know the tangent line: y-yo = m(x-xo) the normal line is just: y-yo = -1/m (x-xo)

OpenStudy (amistre64):

7 e^7y = 7 e^7y = 1 y = 0

OpenStudy (amistre64):

well that neatens this up alot lol

OpenStudy (amistre64):

e^(7y) (49y'+7y +1) = 0 ; y=0 e^(0) (49y'+0 +1) = 0 49y'+1 = 0 y' = -1/49

OpenStudy (amistre64):

y' is the slope of the tangent; therefore the perp to it is just 49

OpenStudy (amistre64):

so, just construct the normal line using a slope of 49 and a given point of (7,0)

OpenStudy (anonymous):

yayy i got it !!! y= 49x-343 ?

OpenStudy (anonymous):

since the negative reciprocal of -1/49 is the slope and is 49 ?

OpenStudy (anonymous):

or no... i read it from somewhere ?

OpenStudy (amistre64):

that looks pretty good to me

OpenStudy (anonymous):

thank youuu!!!

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=xe%5Exy+%3D+7%2C+y%3D+49x-343%2C+y%3D-%28x-7%29%2F49 the picture it shows is not very good, but it looks like it all works out ;) good luck